### v2

# Noise-induced synchronization in coupled quantum oscillators

Abstract
We consider the quantum dynamics of a pair of coupled quantum oscillators coupled to a common correlated dissipative environment. The resulting equations of motion for both the operator moments and covariances can be integrated analytically using the Lyapunov equations. We find that for fully correlated and fully anti-correlated environments, the oscillators relax into a phase-synchronized state that persists for long-times when the two oscillators are nearly resonant and (essentially) forever if the two oscillators are in resonance. This can be traced to the symmetry of the Lindblad dissipator can lead to strong damping in one region of state space and underdamping in others. In the extreme cases of fully correlated or fully anti-correlated environments, specific regions of state space are fully decoupled from the environment. We also show that the environmental noise correlation leads to quantum entanglement, and all the correlations between the two oscillators are purely quantum mechanical in origin. This work provides a robust mathematical foundation for understanding how long-lived exciton coherences can be linked to vibronic correlation effects.
I Introduction
While recovering from an illness in 1665, Christiaan Huygens observed that two identical pendulum clocks, when suspended from a common substantial beam, quickly synchronized with the same periods and amplitudes. [1, 2, 3, 4, 5, 6, 7, 8]. He described this intriguing behavior as ”an odd sympathy” and concluded that synchronization occurs due to the small vibrations transmitted through the support, which acts as a coupling mechanism between the pendula. Over time, these interactions facilitate an exchange of energy that synchronizes their motion. He presented these results to the Royal Society using a series of complex geometric arguments to explain the synchronization mechanism[8, 2].
The concept of synchronization has wide-ranging implications across various scientific fields, including the coordination of biological rhythms, the operation of power grids, and the development of synchronized communication systems [9, 10, 11, 12, 13, 14]. On the microscopic level, synchronization is crucial for in terms of developing quantum technologies that are robust against environmental effects such as decoherence, as well as in understanding the underlying mechanisms for the transport of quantum excitations in photosynthetic systems in which observed long-lived exciton coherences are linked to vibronic correlation effects. The enhanced excitonic coherence has a profound influence on the enhancement of exciton transport and delocalization [15, 16, 17, 18, 19, 20].
[…]
We recently proposed that synchronization between pairs of qubits could be enhanced by coupling to a noisy ancillary environment with correlated or anti-correlated coupling.[21, 22] In these works, we show that noise correlation can help to preserve the fidelity and purity of a prepared state and determines the degree of phase synchronization such that anti-correlated noise leads to anti-phase synchronization and correlated noise leads to in-phase synchronization. In this work, we consider a pair of coupled quantum oscillators coupled to a common correlated dissipative environment. We show that correlations within the environment act as a shared resource that leads to entanglement in the steady-state solutions. The model is applicable to a wide range of physical systems from coherent exciton transport in light-harvesting complexes to quantum phase transitions in driven systems. Hence, we conclude that spontaneous synchronization is a universal hallmark of the interaction a quantum system with a correlated environment.
II Lyapunov equations
We begin by writing a suitable Hamiltonian for this system
$H/\hbar=\omega_{1}a_{1}^{\dagger}a_{1}+\omega_{2}a_{2}^{\dagger}a_{2}+g(a_{1}^{\dagger}a_{2}+a_{2}^{\dagger}a_{1})$ (1)
in which $[a_{i},a_{j}^{\dagger}]=\delta_{ij}$ are bosonic operators with harmonic frequencies $\{\omega_{i}\}$ coupled via $g$ which allows the exchange of quanta from one local mode to the other. For the isolated system, the total number of excitations in the conserved since $[a_{1}^{\dagger}a_{1}+a_{2}^{\dagger}a_{2},(a_{1}^{\dagger}a_{2}+a_{2}^{\dagger}a_{1})]=0$. Further, we can remove the coupling term by transforming to a normal-mode representation without loss of generality. However, for purposes of our analysis we shall continue to work in a local-mode representation. The equations of motion take the general form
$\frac{d\rho}{dt}=-i[H,\rho]+D_{L}(\rho)$ (2)
where $D_{L}(\rho)$ is the Lindblad dissipator which has the general form
$D_{L}(\rho)=L\rho L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,\rho\}$ (3)
where $L$ is a Lindblad operator describing the system/bath interactions.  [^1] We note that our analysis was done within Mathematica (v14.x) by defining a series bosonic algebra rules to generate and simplify the rather cumbersome expressions for multi-component boson systems. Included in this is a similar set of rules and transformations useful for multi-components spin systems. We include a link to a notebook with our derivations in the Data Availability section.
Following from our previous work, and assuming that the coupling to a common environment can be described by correlated Wiener processes, the correlation between the dissipative channels can be introduced by defining a correlation matrix, $\Xi$ which takes the form
$\Xi=\left(\begin{array}[]{cc}1&\xi\\ \xi&1\end{array}\right)$ (6)
in which $-1\leq\xi\leq 1$ serves as a correlation parameter. We use the eigenvalues and corresponding eigenvectors of this to construct new Lindblad operators corresponding to the correlated and anticorrelated components
$L_{\pm}=\sqrt{\frac{1\pm\xi}{2}}(L_{1}\pm L_{2})$ (7)
where $L_{1}$ is a Lindblad operator for oscillator 1 and $L_{2}$ is a Lindblad operator acting on oscillator 2. The total dissipator is then the sum of the symmetric and antisymmetric combinations. When $\xi=0$, the cross components vanish and we have uncorrelated oscillators. When $\xi=1$, the anti-symmetric contribution vanishes leaving only symmetric terms. This implies that symmetric normal modes will be more strongly damped by increasingly correlated environments, essentially leaving the anti-symmetric modes under damped to the extent that for fully correlated environments ($\xi\to 1$), symmetric normal modes of the system will be completely undamped. Likewise, in case of anti-correlated noise ($\xi=-1$), anti-symmetric normal modes are suppressed leading to in-phase synchronization.
In general, the correlation matrix $\Xi$ describes the topological adjacency between correlated Wiener processes such that $dW_{i}(t)dW_{j}(t^{\prime})=\Xi_{ij}\delta(t-t^{\prime})dt$ and may represent a connected graph belonging to a particular discrete symmetry group. Consequently, the eigenvectors of $\Xi$ can be classified according to the irreducible representations of that group.
Since our equations of motion generate a Gaussian map, we will instead work with the operator expectation values and covariances rather than with the density matrix itself. To establish our notation, we write the vector of moments of the system operators as ${\bf x}=\{\langle a_{1}\rangle,\langle a^{\dagger}_{1}\rangle,\langle a_{2}\rangle\langle a_{2}^{\dagger}\rangle\}$ and the covariances as
$\Theta_{ij}=\frac{1}{2}\langle\{\delta x_{i}^{\dagger},\delta x_{j}\}\rangle=\frac{1}{2}\langle\{x_{i}^{\dagger},x_{j}\}\rangle-\langle x^{\dagger}_{i}\rangle\langle x_{j}\rangle$ (8)
It is straightforward to show that $\bf x$ and $\Theta$ evolve according to the following set of equations of motion
$\frac{d{\bf x}}{dt}=W\cdot{\bf x}+f$ (9)
where $W$ is the dynamical matrix and $f$ includes any source or sink terms. Similarly,
$\frac{d\Theta}{dt}=W\cdot\Theta+\Theta\cdot W^{\dagger}+F$ (10)
where matrix $F$ depends only on the dissipative terms in the Liouville equation. These are referred to as the Lyapunov equations and are useful for understanding the stability and dynamics of chaotic systems[24].  [^2] Since these are equations of motion for the expectation values rather than the operators themselves, we use the adjoint dissipator $\langle\hat{O}D(\rho)\rangle={\rm tr}\left[{\hat{O}}\left(L\rho L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,\rho\}\right)\right]=\langle L^{\dagger}\hat{O}L-\frac{1}{2}\{L^{\dagger}L,\hat{O}\}\rangle$ (11) in deriving these terms. For the coupled oscillators, we find the dynamical matrix is given by
$W=\left(\begin{array}[]{cccc}-i\omega_{1}&0&-ig-\xi\gamma_{12}/2&0\\ 0&+i\omega_{1}&0&ig-\xi\gamma_{12}/2\\ -ig-\xi\gamma_{12}/2&0&i\omega_{2}&0\\ 0&ig-\xi\gamma_{12}/2&0&+i\omega_{2}\end{array}\right)-\frac{\gamma}{2}I_{4}$ (16)
where $I_{4}$ is the $4\times 4$ identity matrix. Similarly,
$F=-\left(\begin{array}[]{cccc}\frac{\gamma}{2}+\gamma n_{1}&0&\xi\gamma\sqrt{n_{1}n_{2}}&0\\ 0&\frac{\gamma}{2}+\gamma n_{1}&0&\xi\gamma\sqrt{n_{1}n_{2}}\\ \xi\gamma\sqrt{n_{1}n_{2}}&0&\frac{\gamma}{2}+\gamma n_{2}&0\\ 0&\xi\gamma\sqrt{n_{1}n_{2}}&0&\frac{\gamma}{2}+\gamma n_{2}\\ \end{array}\right)$ (21)
Both $W$ and $F$ can be rearranged into block-diagonal form and from the Lyuapunov equations, we obtain the equations of motion for the moments and covariances.
Starting with the moment equations
$\frac{\partial}{\partial t}\left(\begin{array}[]{c}\langle a_{1}\rangle\\ \langle a_{2}\rangle\end{array}\right)=-i\left(\begin{array}[]{cc}\omega_{1}-i\gamma/2&g-i\xi\gamma_{12}/2\\ g-i\xi\gamma_{12}/2&\omega_{2}-i\gamma/2\end{array}\right)\left(\begin{array}[]{c}\langle a_{1}\rangle\\ \langle a_{2}\rangle\end{array}\right)$ (28)
The non-Hermitian dynamical matrix has complex eigenvalues
$\lambda_{\pm}=\frac{1}{2}(\omega_{1}+\omega_{2}-i\gamma)\pm\frac{1}{2}\sqrt{(2g-i\xi\gamma_{12})^{2}+(\omega_{1}-\omega_{2})^{2}}$ (29)
where
$\gamma_{12}$ $=\gamma(\sqrt{(n_{1}+1)(n_{2}+1)}-\sqrt{n_{1}n_{2}})$ (30)
is a temperature dependent relaxation rate. If we take the case where the two oscillators are decoupled ($g=0$), synchronization becomes spontaneous when $\gamma_{12}\xi>(\omega_{1}-\omega_{2}$. Below the synchronization regime, both symmetric and anti-symmetric normal modes are equally damped at a rate given by $\gamma/2$. However, within the synchronization regime one normal mode becomes strongly damped while the other mode is more weakly damped. Only when the two modes are in resonance with $\omega_{1}=\omega_{2}$ does synchronization occur for all values of $\xi$. Similar regimes have been suggested by a series of recent papers concerning harmonic chains coupled by dissipative nearest neighbor terms [26, 27, 28, 29].
Figure  1(a,b) shows the real and imaginary eigenvalues of the dynamical matrix in Eq. 28 for a $T=0K$ system with detuning $\delta=$ and relaxation rate $\lambda=$. Here, we set $\omega_{1}=1$ and $\omega_{2}=\omega_{1}+\delta$ to set the time scale of our system. The imaginary terms are also offset by $\lambda/2$. Below a critical value of $\xi$, the oscillators are damped as usual. We refer to this as the normal regime. At $\xi_{crit}=(\delta/\gamma_{12})$, we have an exceptional point where the two real normal mode frequencies coalesce, and one mode becomes less damped while the other becomes more strongly damped. As $\xi\to\pm 1$, one of the two normal modes will become completely decoupled from the dissipative environment.
Figure 1: Real (a) and Imaginary (b) components of the eigenvalues of the non-Hermitian dynamical matrix in Eq. 28 for various parametric values: $\delta=\omega_{1}-\omega_{2}=0.2$, $\gamma=0.5$ and increasing values of exchange coupling $g=0$ to $g=0.04$ The system enters the spontaneous synchronized domain above the critical coupling $\xi_{crit}=\delta/\gamma$
In Fig. 2 we show some representative trajectories for $\langle a_{1}(t)\rangle$ and $\langle a_{2}(t)\rangle$ for the case of zero detuning ($\delta=0$) and with various degrees of environmental correlations. We see clearly that anti-correlated noise ($\xi=-1$) leads to in-phase synchronization while fully correlated noise ($\xi=+1$) leads to anti-phase synchronization. Further more, at long time, the synchronized states are fully decoupled from the environment and do not decay as compared to the fully uncorrelated case which relaxes to zero on the time-scale dictated by $\gamma_{12}$.
Figure 2: Representative dynamics for systems driven by anti-correlated (a), uncorrelated (b), and correlated (c) noise. (Parameters: $\delta=0$, $g=0.1$, $\gamma=0.1$)
We now turn our attention towards the equations of motion for the moments as given by Eq.10. The time-evolution is now entirely specified by the moments and covariance matrix. Representative trajectories for the moments are given in Fig.1(a-c). Here, we initiate the dynamics from the same initial condition (corresponding to to two displaced oscillators), however, with differing degrees of environmental correlation. In Fig. 1(a) and (c), phase synchronization or anti-synchronization is achieved after only a few oscillation periods, after which the dynamics settles into a steady-state limit cycle. In contrast, for uncorrelated environments, the two essentially independent oscillators relax as one would expect.
II.1 Steady State Solutions
It is useful to examine the long-time or stationary state solutions in which
$W\cdot\tilde{\Theta}+\tilde{\Theta}\cdot W^{\dagger}=-F.$ (31)
The Lyapunov equation has a closed form solution for $\tilde{\Theta}$ when $\omega_{1}=\omega_{2}$ [30],
$\tilde{\Theta}_{11}$ $=\langle a_{1}^{\dagger}a_{1}\rangle_{ss}+1/2=(\overline{n}_{1}+\frac{1}{2})+\frac{2g^{2}}{4g^{2}+\gamma^{2}}(\overline{n}_{2}-\overline{n}_{1})+\xi^{2}\gamma_{12}\Delta$ (32)
$\tilde{\Theta}_{22}$ $=\langle a_{2}^{\dagger}a_{2}\rangle_{ss}+1/2=(\overline{n}_{2}+\frac{1}{2})-\frac{2g^{2}}{4g^{2}+\gamma^{2}}(\overline{n}_{2}-\overline{n}_{1})+\xi^{2}\gamma_{12}\Delta$ (33)
$\tilde{\Theta}_{12}$ $=\langle a_{1}^{\dagger}a_{2}\rangle_{ss}=i\frac{g\gamma}{4g^{2}+\gamma^{2}}(\overline{n}_{2}-\overline{n}_{1})+\gamma\xi\Delta$ (34)
where $\Delta$ is a collection of variables given by
$\Delta=\frac{\left(\gamma_{12}\left(n_{1}+n_{2}+1\right)+2\gamma\sqrt{n_{1}n_{2}}\right)}{2\left(\gamma^{2}-\gamma_{12}^{2}\xi^{2}\right)}.$ (35)
This term imposes an additional constraint on the range of $\xi$ since it becomes singular when $\overline{n}_{1}\neq\overline{n_{2}}$
$\xi=\pm\sqrt{\frac{1}{\overline{n}_{1}+\overline{n}_{2}+1}},$ (36)
otherwise, $-1\leq\xi\leq 1$. In what follows, we focus our attention on the special case where the two oscillators are at the same frequency but may be held at different temperatures.
The off-diagonal $\tilde{\Theta}_{12}$ terms corresponds to the steady-state “flow” of quanta between the two oscillators. Interestingly, if there is a temperature bias between the two oscillators, we develop correlation since $\tilde{\Theta}_{12}\neq 0$. In fact, if we consider the time evolution of the population of oscillator 1:
$\frac{dn_{1}(t)}{dt}=\gamma(\overline{n}_{1}-n_{1}(t))-ig(\langle a_{1}^{\dagger}a_{2}\rangle+\langle a_{2}^{\dagger}a_{1}\rangle).$ (37)
This is a continuity equation since we have quanta entering from the bath (1st term) and leaving to the other oscillator (2nd term). For the stationary state, the two must be equal
$\gamma(\overline{n}_{1}-n_{1}(t))=ig(\langle a_{1}^{\dagger}a_{2}\rangle+\langle a_{2}^{\dagger}a_{1}\rangle)=J$ (38)
and
$J=\frac{2g\gamma}{4g^{2}+\gamma^{2}}(\overline{n}_{2}-\overline{n}_{1})$ (39)
gives the flux of quanta from the warmer to cooler baths. This does not depend upon correlations within the environment and only depends upon the exchange coupling term within the Hamiltonian itself.
The covariance matrix, $\Theta$, must be positive, semi-definite and should satisfy the uncertainty principle at all times. This can be stated compactly as
$\Theta+i\Omega\geq 0$ (40)
where $\Omega$ is the symplectic form
$\Omega=\oplus_{n=1}^{N}\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right).$ (43)
For the case at hand, we can see that eigenvalues of $\Theta$ are always positive, semi-definite. Likewise, the symplectic form remains invariant under any symplectic transformation, including the one that diagonalizes $\Theta$ (Williamson’s Theorem)[31]. Consequently, $\Theta$ represents a bona fide covariance matrix for a physical gaussian state [32].
II.2 Information Theory for Gaussian States
Figure 3: (a) Mutual information shared by two oscillators in the presence of a temperature bias as specified by the average occupations $n_{1}$ and $n_{2}$. (b) Mutual information shared by two oscillators at a common temperature. In each case, we set $\omega_{1}=\omega_{2}=1$, $g=1$, and $\gamma=0.1$. The vertical dashed lines correspond to the limiting conditions imposed by Eq.36 when $n_{1}\neq n_{2}$.
The total entropy of a bi-partite system is strongly sub-additive in that $S_{A}+S_{B}\geq S_{AB}$. This holds true for the von Neumann entropy but is not true in general for the various orders of the Rényi entropies. However, strong sub-additivity for Rényi order 2 does hold for Gaussian states and is simply related to the purity[32]
$S_{2}=-\ln{{\rm tr}(\rho^{2})}$ ) (44)
One can easily show that for the 2-mode case,
$S_{2}(\Theta)=\frac{1}{2}\ln{{\rm det}\Theta}+N\ln{2}.$ ) . (45)
Furthermore, since the joint covariance matrix can be written in block form as
$\Theta_{AB}=\left[\begin{array}[]{cc}\Theta_{A}&\Gamma\\ \Gamma^{\dagger}&\Theta_{B}\end{array}\right]$ (48)
then the covariance within subspace A or B is given by the two diagonal blocks and the correlations between the two are the off-diagonal blocks ($\Gamma$). The mutual information is then simply
$)\displaystyle I_{2}(A:B)=S_{A}+S_{B}-S_{AB}=\frac{1}{2}\ln{\frac{\text{det}\Theta_{A}\text{det}\Theta_{B}}{\text{det}\Theta_{AB}}}$ (49)
where we used the Rényi-2 entropy. This mutual information is a measure of the total correlation between the two oscillators.
Fig. 3a shows steady-state mutual information, $I_{2}(A:B)$, for various cases where there is a thermal bias between the two oscillators. $I_{2}(A:B)$ diverges as $\xi$ approaches the limit imposed by Eq. 36. Since the fully synchronized (or anti-synchronized) steady-state is only achieved when $\xi=\pm 1$, a thermal bias will inhibit the formation of this state. Fig.3(b) illustrates the case where the two oscillators are at a common temperature. Here, $I_{2}(A:B)$ diverges at $\xi=\pm 1$. This means that in the fully synchronized limit, a measurement on subsystem $A$ gives you full access to the information in subsystem $B$.
Discord
For the given bi-partite system and the specific form of the covariance matrix, it is possible to filter out the classical part of the total mutual information as we showed in our previous work [22]. Following the work by Adesso, Datta et al.[32, 33] we can define the measurement-dependent classical part of the total mutual information by $\mathcal{J}_{2}(\rho_{A|B})\geq 0$. This leaves us with a lack of information about subsystem A when a Gaussian measurement is performed on subsystem B. Such measurements preserve the Gaussian nature of states. We have
${\cal J}_{2}(\rho_{A|B})=\sup_{\Sigma^{\Pi}_{B}}\frac{1}{2}\ln\left(\frac{\text{det}\Theta_{A}}{\text{det}\tilde{\Theta}_{A}^{\Pi}}\right)$ (50)
where
$\tilde{\Theta}_{A}^{\Pi}=\Theta_{A}-\Gamma_{AB}\left(\Theta_{B}+\Sigma^{\Pi}_{B}\right)^{-1}\Gamma^{T}_{AB}$ (51)
is the Shur complement of the $\Theta_{A}$ block and $\Sigma_{B}^{\Pi}$ is the seed of a Gaussian measurement on B. Naturally then the Discord[34] or the degree of quantumness can be defined as the difference between the total correlation and the purely classical contributions[34].
$\begin{split}{\cal D}_{2}&=I_{2}(A:B)-{\cal J}_{2}(A|B)\\ &=\inf_{\Sigma_{B}^{\Pi}}\frac{1}{2}\ln\left(\frac{\text{det}\Theta_{B}\text{ det}\tilde{\Theta}_{A}^{\Pi}}{\text{det}\Theta_{AB}}\right).\end{split}$ (52)
In Ref. 22 we showed that one can obtain a tight lower bound on the amount of quantum correlation between subsystems by simply removing the coherences between the two subspaces within the total density matrix and computing the mutual information between the two remaining diagonal blocks. For Gaussian systems, this follows from Eq. 50 since ${\cal J}_{2}(\rho_{A|B})$ is extremized under this condition and Eq. 52 becomes equal to the total correlation.
In our previous work on coupled qubit dimers, we found that depending upon the correlation within the environment as specified by $\xi$, by and large, coupling to a common reservoir introduces quantum correlations between the spins.[21, 22] For the case at hand, we have two possible sources of correlation in our model. First, the direct coupling introduces correlation when the two oscillators are at different temperatures. Secondly, even when the direct coupling is turned off or the two oscillators are at the same temperature, correlations within the common environmental mode introduces correlations between the two oscillators. Our results are consistent with the formal results in Ref. 32 which show that the only two-mode Gaussian states with zero Gaussian quantum discord are product states, i.e. states with no correlations at all, that constitute a zero measure set. Consequently, all correlated two-mode Gaussian states have nonclassical correlations certified by a nonzero quantum discord.
III Discussion
We demonstrate here that synchronization and entanglement is a universal phenomina that can arise via indirect coupling to a correlated dissipative environmental mode. Environmental correlation imposes a symmetry on the system by selectively increasing the relaxation rates for modes that have the same symmetry as the environmental correlations leaving the remaining modes effectively decoupled from the environment. This selectivity will persist in multi-mode systems, especially in system with a high degree of topological symmetery. One can use this to engineer dissipation and decoherence free domains within the total Hilbert space of the system.
Our results also have implications in terms of understanding the underlying mechanisms for the transport of quantum excitations in photosynthetic systems in which observed long-lived exciton coherences are linked to vibronic correlation effects. The enhanced excitonic coherence has a profound influence on the enhancement of exciton transport and delocalization [15, 16, 17, 18, 19, 20].
One of the biggest challenges in quantum computing is decoherence—the loss of quantum coherence due to interactions with the environment. Quantum states, particularly superpositions and entanglement, are very fragile and prone to disruption due to environmental fluctuations. This work suggests that by carefully engineering the local environment about each qubit and correlations between qubits, one can construct symmetry-adapted baths that force the system into phase-locked state. For instance, synchronized quantum oscillators may continue to operate in coherence for longer durations, increasing the reliability of quantum computations and preserving the integrity of qubits. By actively mitigating decoherence, facilitating error correction, and improving qubit control, synchronization could address several of the most pressing challenges in quantum computing. Additionally, it enables more efficient quantum information transfer, helps to stabilize entanglement by preserving fidelity[21], and supports the development of precise quantum clocks. These advancements will be pivotal in realizing robust and scalable quantum computers.
Acknowledgments
This work was funded by the National Science Foundation (CHE-2404788) and the Robert A. Welch Foundation (E-1337).
Author contribution statement
The authors acknowledge discussions with Prof. Carlos Silva (U. Montreal) and Andrei Piryatinski (LANL) that led to the development of this model and series of calculations. Both authors discussed the results and contributed to the final manuscript.
Data Availability Statement
Data files and source codes used in this work are freely and publicly available on the Wolfram cloud at the URL: https://www.wolframcloud.com/obj/a7bd6168-46c4-47e6-9f01-1755e79553f6
Competing Interests Statement
The authors have no competing interests to declare.